A properly exposed photograph is taken at f/32 and 1/1000 s. What lens f-stop gives the same exposure with a shutter speed of 1/150 s?
Answer #1

if you were shooting on a 100mm lens, a shutter speed of 1/100 of a second would be optimal. Camera blur can be eliminated at these speeds, so at the beginning the focal length is 1000 mm. From this data we can find the aperture by:

\frac{1000\;mm}{x } = 32 \Rightarrow x = \frac{1000\;mm}{32} = 31.25\;mm

therefore, our new f-stop number is:

\frac{150\;mm}{ 31.25\;mm} = 4.8 \Rightarrow \boxed{f-stop = \frac f{4.8}}